OpenStudy (anonymous):

[6.02] Solve the following system. y = x + 3 3x + y = 19 (7, 4) (−4, 7) (4, 7) (4, −7)

OpenStudy (anonymous):

Help me please

terenzreignz (terenzreignz):

There's the easy way to do this, which takes advantage of the fact that there are choices (only recommended if you're in a hurry) and the normal way to do this, which is to solve the system while somehow, pretending there weren't any choices. Which way do you want to go?

OpenStudy (anonymous):

what is one for the hurry???

terenzreignz (terenzreignz):

Well, you just have to test each one of the choices and see which one works.

OpenStudy (anonymous):

how do you do that??? sorry i am not good with math

terenzreignz (terenzreignz):

Here, an example: Let's test the first choice, (7,4) You replace the x with 7 and the y with 4. for the first equation: y = x + 3 replacing.... 4 = 7 + 3 4 = 10 Which is not true. Therefore, the answer is NOT the first choice (7,4) Try this with the other choices. Remember, it has to work for BOTH of your equations, not just one.